In an effort to confirm the results of Michelson experiment in the year 1882 which measured the speed of light, a scientist currently reports her results of 4 trials having a mean of 756 and a standard deviation of 108. (The units she used is not important here.) Find the 95% confidence interval for the true mean of the speed of light from these statistics. Label Conf. Interval;round results to 1 decimal place. Include drawings in your work.
Answer #1

Solution :

Given that,

t /2,df = 3.181

Margin of error = E = t/2,df * (s /n)

= 3.181 * (108 / 4)

Margin of error = E = 171.8

The 95% confidence interval estimate of the population mean is,

- E < < + E

756 - 171.8 < < 756 + 171.8

584.2 < < 927.8

(584.2 , 927.8)